OpenStudy (anonymous):

simplify square root of 5 times square root of 500. Please help will give medals and will fan!!!!

OpenStudy (anonymous):

\[\sqrt{5}*\sqrt{500}\]

OpenStudy (anonymous):

44

OpenStudy (anonymous):

what is 5 times 500?

OpenStudy (anonymous):

2500

OpenStudy (kc_kennylau):

What's its square root?

OpenStudy (anonymous):

yes what is the square root of 2500?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

The square root of 5 is approximately 2, and the square root of 500 is approximately 22, and 2 x 22 = 44.

OpenStudy (anonymous):

that is the answer!

OpenStudy (kc_kennylau):

@AnswerMyQuestions don't freaking approximate things when you don't freaking have to

OpenStudy (anonymous):

thank you @hello1213

OpenStudy (anonymous):

Who said you don't?

OpenStudy (kc_kennylau):

I say so

OpenStudy (anonymous):

hello can you help me more plz

OpenStudy (anonymous):

You're welcome!

OpenStudy (kc_kennylau):

One formula: \(\sqrt{ab}=\sqrt a\sqrt b\)

OpenStudy (anonymous):

0.125

OpenStudy (anonymous):

let's give @kc_kennylau a turn!

OpenStudy (kc_kennylau):

And \(\sqrt{\frac ab}=\frac{\sqrt a}{\sqrt b}\)

OpenStudy (anonymous):

\[\sqrt{8/18}\] sorry wrote it wrong

OpenStudy (anonymous):

simplify

OpenStudy (kc_kennylau):

Simplify the fraction inside first

OpenStudy (anonymous):

1

OpenStudy (kc_kennylau):

Simplify 8/18 first

OpenStudy (kc_kennylau):

Notice that both 8 and 18 are even

OpenStudy (anonymous):

4/9

OpenStudy (kc_kennylau):

Now apply square root to both the numerator (4) and the denominator (9)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[1/\sqrt{8}\]

OpenStudy (anonymous):

A. 2/3 B. 2 C.-2/3 D.3/2

OpenStudy (kc_kennylau):

What is the square root of 4? And that of 9?

OpenStudy (anonymous):

square root of 4 is 2 and of 9 is 3

OpenStudy (kc_kennylau):

Therefore the square root of 4/9 is...

OpenStudy (anonymous):

A. 2/3 B. 2 C.-2/3 D.3/2

OpenStudy (kc_kennylau):

I'm asking you...

OpenStudy (kc_kennylau):

\[\sqrt{\frac49}=\frac{\sqrt4}{\sqrt9}=?\]